# HiReview: Hierarchical Taxonomy-Driven Automatic Literature Review Generation

### v1

@@ 5.2 Evaluation
Table 1: Results of literature review generation by pure LLMs, naive RAG-based LLMs, AutoSurvey and HiReview. The best performance is in Bold. LLM indicates the LLM is provided top-500 relevant papers retrieved by naive BM25$\uparrow$ indicates that a higher metric value corresponds to better model performance.
@@
Main Results. As shown in Table 1, Our method HiReview consistently outperforms the other review generation methods in all metrics. It excels across all LLMScore categories, with notably high structure (0.9484) and relevance (0.9428) scores. AutoSurvey employs a structured methodology that combines retrieval, outline generation, and section drafting, leading to superior content generation compared to naive systems (with average LLMScore of 0.8957).
Pure LLMs and naive RAG-based LLMs struggle with both stability and performance, which makes them unreliable for consistent literature review generation. AutoSurvey reduces this instability through prompt design and multi-output generation, achieving Structure ±0.05 and Relevance ±0.04—lower deviations than those of pure and naive RAG-based LLMs. HiReview, however, outperforms all other models across all metrics, with consistently low standard deviations. This demonstrates HiReview’s superior stability and consistency in generating high-quality reviews (RQ3). Its success can be attributed not only to HiReview’s use of a graph-context-aware retrieval method but also to the taxonomy tree, which provides hierarchical context for domain-specific concerns within the large language model. A detailed analysis contrasting the outline-then-generation with the taxonomy-then-generation, based on a specific generation example, is provided in Appendix A.5. Furthermore, an example of the generation for a cluster with the central topic Continual Text Classification is included in Appendix A.7.
5.3 Ablation Study
@@
Table 2: Ablation study results for HiReview.
@@
[…]
Impacts of Components. As shown in Table 2, we test three variants of HiReiview mode. HiReview w/o retrieval refers to the variant where the graph retrieval module is removed, and all papers in the citation network are used. A significant drop is observed across all metrics, particularly in Coverage (0.9163 $\rightarrow$ 0.6705) and Relevance (0.9428 $\rightarrow$ 0.7073). This indicates that the inclusion of unrelated papers introduces substantial noise, negatively impacting both the taxonomy tree generation (due to an excess of negative samples in hierarchical clustering) and content generation (where the noise hinders the creation of precise summaries). As a result, the quality opf generated summaries is even worse than those produced by zero-shot LLMs.
HiReview w/o clustering bypasses the clustering process and directly uses the retrieved papers to generate the taxonomy. Instead of iteratively generating topics at each level, this variant creates the taxonomy in a single step. It is marked with * because the topic generator in this case is an LLM i.e., GPT-4o, rather than a fine-tuned LLaMA, as the number of taxonomy trees is insufficient for effective fine-tuning. Although this variant performs worse than HiReview, it still delivers competitive performance, outperforming naive RAG-based LLMs and AutoSurvey. This suggests that the combination of the graph retrieval module and the taxonomy-then-generation paradigm is more effective than naive retrieval-then-generation and outline-then-generation approaches.
HiReview w/o taxonomy removes the hierarchical taxonomy tree generation module, instead using paper clusters to prompt the LLM for both topic generation and content creation. The absence of a hierarchical taxonomy reduces the model’s ability to leverage topic relations across different levels, leading to less organized and relevant content (Structure: 0.9484 $\rightarrow$ 0.8790 and Coverage: 0.9163 $\rightarrow$ 0.8612). Similar to HiReview w/o clustering, HiReview w/o taxonomy does not use fine-tuned LLaMA, and its performance is more degraded. This indicates that when using pure LLM generation methods, generating a hierarchical taxonomy tree to guide content generation significantly enhances the quality of the output. Finally, we can answer remaining questions raised at the beginning of the Experiment Section.
Table 3: Accuracy of hierarchical clustering.
@@
RQ1. We consider two baseline methods for the hierarchical clustering module: one that utilizes an LLM (i.e., GPT-4o) to cluster papers and another that applies $K$-means, adjusting the number of clusters to represent different levels. However, neither method can be jointly trained with the topic generator. Even disregarding the training requirement, the hierarchical nature of the literature review’s taxonomy tree requires soft clustering at the initial layer and hard clustering at subsequent layers—an issue that no existing work addresses. As shown in Table 3, when considering the clustering task alone, both baselines underperform compared to our hierarchical approach. Although the LLM is prompted to perform soft clustering at the first level, offering a slight improvement over $K$-means, it still does not achieve the effectiveness of our hierarchical clustering approach.
@@
RQ2. As shown in Table 1, HiReview, which incorporates a taxonomy tree, outperforms all other models, particularly in Structure, achieving a score of 0.9484. In contrast, AutoSurvey, which follows an outline-then-generation approach without hierarchical taxonomy, shows lower scores, such as a Structure score of 0.9122. The ablation study further supports this. When the taxonomy is removed, the structure score drops significantly (as in Table 2). This demonstrates that the taxonomy tree plays a critical role in organizing and guiding the content generation process, especially when maintaining a clear structure is crucial for a literature review. The taxonomy ensures more coherent and relevant summaries. Without providing the taxonomy tree, the generation loses its hierarchical guidance, leading to less structured and less comprehensive content.
6 Conclusion
In this paper, we propose HiReview, a novel method for literature review generation that leverages a hierarchical taxonomy-driven approach, integrating graph-based clustering with graph context-aware retrieval-augmented large language models. Extensive experiments show that HiReview consistently outperforms state-of-the-art methods across multiple metrics. The ablation study provides strong evidence of the critical role that the hierarchical taxonomy tree plays in guiding the content generation process, leading to more coherent and comprehensive reviews. Compared to the outline-then-generation framework, the taxonomy-then-generation approach offers improved robustness and performance. HiReview achieves a more structured and contextually aligned synthesis of scientific literature than traditional outline-based methods. This work addresses key challenges in automatic literature review generation, particularly in managing large citation networks and preserving the structural integrity of scientific literature reviews. The results establish a new benchmark in the field and emphasize the importance of incorporating both structural and content-based guidance in LLM-driven review generation.
@@ A.1 Dataset
Citation Network Construction Process. For each literature review, we first extracted its references and constructed a citation tree, with the review paper as the root and its cited papers as the leaves. We then repeated this process for each cited paper, constructing a citation tree for each one. Next, we merged all these trees into a single, large citation network, consolidating any duplicate nodes. To automate this process, we used citation information from arXiv, which provides the LaTeX source code for each paper, including the bib or bbl files. If a paper was available on arXiv, we extracted its .tex file to obtain both the abstract and full text, using these as high-quality text features for the corresponding node. We used the arXiv API to automate this process. For papers not available on arXiv, we used the Google Scholar API to automatically retrieve the abstract, which we used as the text feature for the corresponding node in the citation network. Finally, we removed the node representing the original literature review, leaving 1-hop, 2-hop, and 3-hop citation networks for each review.
A.2 Implementation
All experiments were conducted on a Linux-based server equipped with 4 NVIDIA A10G GPUs. For 518 review papers, we successfully collected taxonomy trees with hierarchical clustering labels for 313 of the literature reviews. Of these, 200 reviews were used to train the hierarchical clustering and taxonomy generation module, while the remaining 118 were used to test the performance of the pre-trained hierarchical clustering model. 318 reviews were reserved for a comprehensive evaluation of review content generation. The number of articles retrieved in retrieval phase was set to 200. The scaling factor $\alpha$ is set to 1.
Pre-Train Hierarchical Clustering Module. GNN used in this paper is GAT (Veličković et al., 2017) which has 2 layers with 4 heads per layer and a hidden dimension size of 1024. MLPϕ has 2 layers and a hidden dimension size of 1024. The edge connection threshold $p_{\tau}$ is searched in [0.1, 0.2, 0.5, 0.8]. The clustering model is trained for a maximum of 500 epochs using an early stop scheme with patience of 10. The learning rate is set to 0.001. The training batch is set to 512 and the test batch is 1024.
@@ A.2 Implementation
A.3 Investigation of Retrieval Models
We experimented with different retrieval models and strategies, testing two representative methods: the sparse retrieval model, BM25 (Robertson et al., 2009), and the dense retrieval model, SentenceBert (Reimers, 2019). In citation networks, neighbor information and the topological structure play a crucial role in retrieval, as papers on the same topic often cite each other. To assess the impact of using neighbor information, we applied two retrieval strategies for both models: one incorporating neighbor information as described in Section 4.1 (Retrieval w/ Neighbor) and the other excluding neighbor information (Retrieval w/o Neighbor). Given a topic (specifically, the title of a review paper), we retrieved papers related to this topic from the citation network and measured the accuracy by calculating how many of the retrieved papers appeared in the references of the corresponding literature review. The number of retrieved papers was not fixed, but matched the reference count for each review paper.
Table 4: Results of retrieval on the citation network corresponding to 50 review papers. 2-hop and 3-hop represent citation networks of review papers at different scales. 1-hop (merged) refers to the 1-hop citation network of a review paper, merged with all other 1-hop citation networks, different review papers. Similarly, 2-hop (merged) is constructed by merging the 2-hop citation network of a review with all other 49 review papers.
@@ A.3 Investigation of Retrieval Models
A.4 The Choice of GNN
In addition to GAT (Veličković et al., 2017), we also explored other GNNs as graph encoders, i.e., GCN (Kipf & Welling, 2016) and Graph Transformer (Shi et al., 2020). The comparison results of these models on clustering are shown in Table 5.
Table 5: Performance of different GNN on hierarchical clustering.
@@ A.4 The Choice of GNN
A.5 Comparison to Outline-Then-Generation
Recently, the paradigm for generating literature reviews has shifted from extractive models to outline-based generative models, where LLMs are used to generate an outline of the literature review and then the outline is used to guide generation of content (i.e., outline-then-generation in (Wang et al., 2024, Wu et al., 2024)). Taking Lifelong Learning of Large Language Models as an example topic, we demonstrate the generation of HiReview using two different paradigms: outline-then-generation and taxonomy-then-generation. We analyze the gains and losses of each approach. For the outline-then-generation paradigm, we allow the LLM to generate the outline based on clusters at the base level.
Comparison. As illustrated in Fig. 4 and Fig. 5, both the human-designed taxonomy and the taxonomy generated by HiReview for Lifelong Learning of LLMs provide a clear and cohesive hierarchical structure. In contrast, the outline in Fig. 3 covers plausible approaches and concepts related to lifelong learning and organizes these elements. The outline provides a broad structure covering everything from basics to future trends. For instance, it includes Introduction and Future Directions and Emerging Trends, allowing for an overview of the literature review. The taxonomy clearly outlines specific applications like Continual Relation Extraction, allowing for a more focused discussion on particular areas of lifelong learning. Therefore, use outline-then-generation for comprehensive, structured reviews that cover both theoretical and practical aspects, making it particularly effective for diverse audiences. In contrast, use taxonomy-then-generation for more focused, task-specific reviews, especially when the emphasis is on core concepts, and the audience is already familiar with the basics of a specific domain.
When to use outline-then-generation? The outline offers a logical and detailed structure that ensures comprehensive coverage of all important aspects of the topic. It follows a clear progression from introduction to conclusion, making it easier for readers to follow. This approach includes broad coverage, addressing theoretical foundations, architectures, training methods, applications, challenges, evaluation methods, and future trends. However, it places less emphasis on specific lifelong learning scenarios, as the outline doesn’t explicitly highlight particular tasks or domains where lifelong learning is applied. The outline-then-generation paradigm is ideal when a structured, logically flowing review is needed, covering the topic comprehensively from foundational principles to advanced applications.
When to use taxonomy-then-generation? The taxonomy clearly highlights different domains and tasks where lifelong learning is applied (e.g., relation extraction, semantic segmentation). It emphasizes core lifelong learning concepts, directly addressing key issues like catastrophic forgetting and lifelong learning strategies. This approach offers flexibility, allowing for easy addition of new categories or tasks as the field evolves. It also provides a concise overview, offering a quick snapshot of the main areas of research in lifelong learning for LLMs. However, it may be less comprehensive, missing broader contexts such as theoretical foundations, evaluation methods, and future directions. Additionally, it doesn’t provide a narrative flow from basic concepts to advanced applications. The taxonomy-then-generation paradigm is best suited for highlighting specific areas, offering flexible categorization, and providing a concise overview of research without requiring a detailed or linear narrative.
{forest} my tree [Lifelong Learning of Large Language Models [1. Introduction to Lifelong Learning of Large Language Models [1.1. Concept of Lifelong Learning] [1.2. Large Language Models: An Overview] [1.3. Importance of Lifelong Learning in LLMs] [1.4. Recent Research Trends] ] [2. Theoretical Foundations of Lifelong Learning in Large Language Models [2.1. Principles of Lifelong Learning] [2.2. Cognitive and Neuroscience Inspirations] [2.3. Mathematical Frameworks for Lifelong Learning] [2.4. Adaptation of Lifelong Learning to LLMs] ] [3. Large Language Model Architectures [3.1. Transformer-based Architectures] [3.2. Mixture of Experts (MoE) Models] [3.3. Retrieval-Augmented Models] [3.4. Multimodal Architectures] ] [4. Efficient and Effective Training Methods for Lifelong Learning Models [4.1. Continual Learning Techniques] [4.2. Instruction Tuning and In-Context Learning] [4.3. Retrieval-Augmented Generation] [4.4. Meta-Learning and Few-Shot Learning] [4.5. Knowledge Distillation and Model Compression] ] [5. Applications and Use Cases of Lifelong Learning Models [5.1. Natural Language Processing Tasks] [5.2. Multilingual and Cross-lingual Applications] [5.3. Code Generation and Understanding] [5.4. Domain-Specific Applications] [5.5. Multimodal Applications] ] [6. Challenges and Limitations [6.1. Catastrophic Forgetting] [6.2. Computational Resources and Efficiency] [6.3. Data Privacy and Security] [6.4. Bias and Fairness Issues] [6.5. Ethical Considerations] ] [7. Evaluation and Benchmarking of Large Language Models [7.1. Performance Metrics for Lifelong Learning] [7.2. Benchmark Datasets and Tasks] [7.3. Evaluation of Generalization and Adaptation] [7.4. Robustness and Consistency Measures] ] [8. Future Directions and Emerging Trends [8.1. Scaling Laws and Model Size] [8.2. Integration with External Knowledge Bases] [8.3. Personalization and Adaptability] [8.4. Interpretability and Explainability] [8.5. Autonomous Learning and Self-Improvement] ] [9. Conclusion [9.1. Summary of Key Insights] [9.2. Open Research Questions] [9.3. Potential Future Research Avenues] ] ] Figure 3: Outline Generated by HiReview.
{forest} my tree [Lifelong Learning of Large Language Models [Continual Pretraining [Continual Vertical Domain Pretraining] [Continual Language Domain Pretraining] [Continual Temporal Domain Pretraining] ] [Continual Finetuning [Continual Text Classification] [Continual Named Entity Recognition] [Continual Relation Extraction] [Continual Machine Translation] [Continual Instruction Tuning] [Continual Knowledge Editing] [Continual Alignment] ] [External Knowledge [Retrieval-Based Lifelong Learning] [Tool-Based Lifelong Learning] ] ] Figure 4: Taxonomy Tree Built by Human (Zheng et al., 2024).
{forest} my tree [Lifelong Learning of Large Language Models [Continual Pre-training of Language Models] [Overcoming Catastrophic Forgetting] [Lifelong Learning Strategies] [Continual Learning in Specific Domains [Continual Relation Extraction] [Continual Semantic Segmentation] [Continual Object Detection] [Continual Learning in Reinforcement Learning] ] [Applications of Continual Learning [Large Language Model Editing and Tuning] [Continual Learning in Task-Oriented Dialogue Systems] [Continual Event Detection] [Continual Text Classification] [Continual Learning for Aspect Sentiment Classification Tasks] ] ] Figure 5: Taxonomy Tree Generated by HiReview.
A.6 Prompt Used
Prompt for generating content of each cluster.
Instruction: You are writing an overall and comprehensive literature review about [TOPIC].   The taxonomy tree of this literature review is:
[OVERALL TAXONOMY TREE]
Now, you need to write the content for a section: [TOPIC OF CLUSTER]. The following is a list of references: [TITLE 1]: [CONTENT 1] [TITLE 2]: [CONTENT 2] [TITLE N]: [CONTENT N]
Requirements:
The subsection should contain more than [WORD NUM] words. When writing sentences based on specific papers, cite the paper using a numbered reference format [X], where X is the number corresponding to the paper in the reference list at the end of the document. The full titles of the papers should be listed in this reference section. Citation Guidelines: Summarizing Research: Cite sources when summarizing existing literature. Using Specific Concepts or Data: Provide citations when discussing specific theories, models, or data. Using Established Methods: Cite the creators of methodologies you employ in your literature review. Supporting Arguments: Cite sources that back up your conclusions and arguments. Only return the content with more than [WORD NUM] words that you write for the section [TOPIC OF CLUSTER] without any other information:
Prompt for evaluating the coverage of generated review.
Instruction: You are an expert in literature review evaluation, tasked with comparing a generated literature review to a human-written literature review on the topic of [TOPIC].   Human-Written Literature Review (Gold Standard):
[GROUND TRUTH REVIEW]
Generated Literature Review (To be evaluated): [GENERATED REVIEW]
Evaluation Requirements: The human-written literature review serves as the gold standard. Your job is to assess how well the generated literature review compares in terms of coverage. Carefully analyze both reviews and provide a score.
Evaluate Coverage (Score out of 100). Assess how comprehensively the generated review covers the content from the human-written review. Consider:
The percentage of key subtopics addressed from the human-written review. The depth of discussion for each subtopic compared to the human-written version. Balance between different areas within the topic as presented in the human-written review. Only return only a numerical score out of 100, where 100 represents perfect alignment with the human-written literature review, without providing any additional information.
Prompt for evaluating the structure of generated review.
Instruction: You are an expert in literature review evaluation, tasked with comparing a generated literature review to a human-written literature review on the topic of [TOPIC].   Human-Written Literature Review (Gold Standard):
[GROUND TRUTH REVIEW]
Generated Literature Review (To be evaluated): [GENERATED REVIEW]
Evaluation Requirements: The human-written literature review serves as the gold standard. Your job is to assess how well the generated literature review compares in terms of structure. Carefully analyze both reviews and provide a score.
Evaluate Structure (Score out of 100). Assess how well the generated literature review’s organization and flow match that of the human-written literature review. Consider: Similarity in logical progression of ideas. Presence of a clear hierarchy of sections and subsections comparable to the human-written literature review. Appropriate use of headings and subheadings in line with the human-written version. Overall coherence within and between sections relative to the human-written literature review. Only return only a numerical score out of 100, where 100 represents perfect alignment with the human-written literature review, without providing any additional information.
Prompt for evaluating the relevance of generated review.
Instruction: You are an expert in literature review evaluation, tasked with comparing a generated literature review to a human-written literature review on the topic of [TOPIC].   Human-Written Literature Review (Gold Standard):
[GROUND TRUTH REVIEW]
Generated Literature Review (To be evaluated): [GENERATED REVIEW]
Evaluation Requirements: The human-written literature review serves as the gold standard. Your job is to assess how well the generated literature review compares in terms of relevance. Carefully analyze both reviews and provide a score.
Evaluate Relevance (Score out of 100). Evaluate how well the generated literature review aligns with the focus and content of the human-written literature review. Consider: Alignment with the core aspects of [TOPIC] as presented in the human-written literature review. Relevance of examples and case studies compared to those in the human-written literature review. Appropriateness for the target audience as demonstrated by the human-written literature review. Exclusion of tangential or unnecessary information not present in the human-written version. Only return only a numerical score out of 100, where 100 represents perfect alignment with the human-written literature review, without providing any additional information.
A.7 Example of Generation

### v2

@@ 5.2.1. Taxonomy Learning Results
Table 1. Results of taxonomy generation. The best performance is in Bold. $\uparrow$ indicates that a higher metric value corresponds to better model performance.
[…]
Large Language Models
GPT-4o $0.9132_{\pm 0.015}$ $0.8914_{\pm 0.018}$ 0.8828±0.012 $0.8958$ 0.6400 2.4400 $0.8376_{\pm 0.014}$
[…]
Taxonomy Learning Methods
HiExpan $0.8427_{\pm 0.021}$ $0.8538_{\pm 0.018}$ $0.8219_{\pm 0.020}$ $0.8395$ 0.4650 1.5000 $0.8145_{\pm 0.015}$
[…]
The potential of taxonomy-guided literature review generation led us to investigation an automatic literature review generation method based on our proposed HiGTL, called Hierarchical Taxonomy-Driven Automatic Literature Review Generation (HiReview). The next section presents the results of this literature review generation.
5.2.2. Review Generation Results
We compare the reviews generated by our model with those written by human experts, zero-shot LLMs and naive RAG-based LLMs, i.e., GPT-4o and Claude-3.5. Zero-shot LLMs rely solely on their pretrained knowledge to generate literature review content, naive RAG-based LLMs utilize a simple BM25 retriever. The LLM backbone of AutoSurvey and HiReview are both GPT-4o for generating the literature review content. Additionally, we benchmark our model against the state-of-the-art review generator, AutoSurvey (Wang et al., 2024). Following AutoSurvey’s scoring criteria, we evaluate the generated reviews based on coverage, structure, and relevance when selecting the best output and calculating the LLMScore. When evaluate the review content, instead of directly scoring the generation, we have LLMs compare the generated content with human-written reviews.
Table 2. Results of literature review generation by pure LLMs, naive RAG-based LLMs, AutoSurvey, and HiReview. The best performance is in Bold. LLM indicates the LLM is provided with the top-500 relevant papers retrieved by BM25. $\uparrow$ indicates that a higher metric value corresponds to better model performance.
@@ 5.2.2. Review Generation Results
As shown in Table 2, Our method HiReview consistently outperforms the other review generation methods in all metrics. It excels across all LLMScore categories, with notably high structure (0.9484) and relevance (0.9428) scores. AutoSurvey employs a structured methodology that combines retrieval, outline generation, and section drafting, leading to superior content generation compared to naive systems (with average LLMScore of 0.8957).
Pure LLMs and naive RAG-based LLMs struggle with both stability and performance, which makes them unreliable for consistent literature review generation. AutoSurvey reduces this instability through prompt design and multi-output generation, achieving Structure ±0.05 and Relevance ±0.04—lower deviations than those of pure and naive RAG-based LLMs. HiReview, however, outperforms all other models across all metrics, with consistently low standard deviations. This demonstrates HiReview’s superior stability and consistency in generating high-quality reviews. Its success can be attributed not only to HiReview’s use of a graph-context-aware retrieval method but also to the taxonomy tree, which provides hierarchical context for domain-specific concerns within the large language model. An example of a generated literature review section is provided in Appendix B.
5.3. Ablation Study
@@
Table 3. Ablation study results for HiReview.
@@
Impacts of Components. As shown in Table 4, we test three variants of HiReiview mode. HiReview w/o retrieval refers to the variant where the graph retrieval module is removed, and all papers in the citation network are used. A significant drop is observed across all metrics, particularly in Coverage (0.9163 $\rightarrow$ 0.6705) and Relevance (0.9428 $\rightarrow$ 0.7073). This indicates that the inclusion of unrelated papers introduces substantial noise, negatively impacting both the taxonomy tree generation (due to an excess of negative samples in hierarchical citation graph clustering) and content generation (where the noise hinders the creation of precise summaries). As a result, the quality opf generated summaries is even worse than those produced by zero-shot LLMs.
HiReview w/o clustering bypasses the clustering process and directly uses the retrieved papers to generate the taxonomy. Instead of iteratively generating topics at each level, this variant creates the taxonomy in a single step. It is marked with * because the taxonomy node verbalizer in this case is an LLM i.e., GPT-4o, rather than a fine-tuned LLaMA, as the number of taxonomy trees is insufficient for effective fine-tuning. Although this variant performs worse than HiReview, it still delivers competitive performance, outperforming naive RAG-based LLMs and AutoSurvey. This suggests that the combination of the graph retrieval module and the taxonomy-then-generation paradigm is more effective than naive retrieval-then-generation and outline-then-generation approaches.
HiReview w/o taxonomy removes the hierarchical taxonomy tree generation module, instead using paper clusters to prompt the LLM for taxonomy node verbalization and review generation. The absence of a hierarchical taxonomy reduces the model’s ability to leverage topic relations across different levels, leading to less organized and relevant content (Structure: 0.9484 $\rightarrow$ 0.8790 and Coverage: 0.9163 $\rightarrow$ 0.8612). Similar to HiReview w/o clustering, HiReview w/o taxonomy does not use fine-tuned LLaMA, and its performance is more degraded. This indicates that when using pure LLM generation methods, generating a hierarchical taxonomy tree to guide content generation significantly enhances the quality of the output. Finally, we can answer remaining questions raised at the beginning of the Experiment Section.
Table 4. Ablation study results for HiReview.
@@
Figure 3. Taxonomy Tree Generated by HiGTL.
Figure 4. Taxonomy Tree Built by Zheng et al..
@@
We consider two baseline methods for the hierarchical citation graph clustering module: one that utilizes an LLM (i.e., GPT-4o) to cluster papers and another that applies $K$-means, adjusting the number of clusters to represent different levels. However, neither method can be jointly trained with the taxonomy node verbalizer. Even disregarding the training requirement, the hierarchical nature of the literature review’s taxonomy tree requires soft clustering at the initial layer and hard clustering at subsequent layers—an issue that no existing work addresses. As shown in Table LABEL:tab:cluster, when considering the clustering task alone, both baselines underperform compared to our hierarchical approach.
@@
As shown in Table 2, HiReview, which incorporates a taxonomy tree, outperforms all other models, particularly in Structure, achieving a score of 0.9484. In contrast, AutoSurvey, which follows an outline-then-generation approach without hierarchical taxonomy, shows lower scores, such as a Structure score of 0.9122. The ablation study further supports this. When the taxonomy is removed, the structure score drops significantly (as in Table 4). This demonstrates that the taxonomy tree plays a critical role in organizing and guiding the content generation process, especially when maintaining a clear structure is crucial for a literature review. The taxonomy ensures more coherent and relevant summaries. Without providing the taxonomy tree, the generation loses its hierarchical guidance, leading to less structured and less comprehensive content.
5.4. Case Study
As shown in Figure 4, the taxonomy generated by HiGTL identifies Continual Text Classification as a key subtopic under Applications of Continual Learning, highlighting its relevance in areas like sentiment analysis, spam detection, and topic categorization. It emphasizes the need for models that can adapt incrementally as new data emerges. While HiGTL captures broad applications effectively, it may lack the finer distinctions found in human-crafted taxonomies (Figure 4), such as domain-specific classifications in legal, medical, or social media contexts. However, this generalized framework ensures semantic coherence and practical usability, providing a solid foundation that human experts can easily refine for tasks like literature review generation.
6. Conclusion
In this paper, we introduced HiGTL, a novel end-to-end framework for automatic taxonomy generation from citation graphs, guided by human-provided instructions or preferred topics. By integrating hierarchical citation graph clustering with taxonomy node verbalization, HiGTL effectively combines the structural relationships inherent in citation networks with the rich textual content of academic papers. Our joint optimization strategy ensures both structural coherence and semantic consistency across hierarchical levels. Extensive experiments on 518 citation graphs from high-quality literature reviews in computer science demonstrated that HiGTL generates meaningful taxonomies. Furthermore, we extended our framework to literature review generation with HiReview, showcasing its practical utility in guiding the creation of coherent, well-structured reviews. Ablation studies confirmed the critical role of hierarchical clustering and taxonomy-guided generation in enhancing performance, while case studies illustrated the flexibility and usability of the generated taxonomies. HiGTL offers a scalable, robust solution for organizing scientific knowledge and supporting downstream tasks such as literature review generation, knowledge discovery, and trend identification.
@@ A.1. Dataset
Citation Network Construction Process. For each literature review, we first extracted its references and constructed a citation tree, with the review paper as the root and its cited papers as the leaves. We then repeated this process for each cited paper, constructing a citation tree for each one. Next, we merged all these trees into a single, large citation network, consolidating any duplicate nodes. To automate this process, we used citation information from arXiv, which provides the LaTeX source code for each paper, including the bib or bbl files. If a paper was available on arXiv, we extracted its .tex file to obtain both the abstract and full text, using these as high-quality text features for the corresponding node. We used the arXiv API to automate this process. For papers not available on arXiv, we used the Google Scholar API to automatically retrieve the abstract, which we used as the text feature for the corresponding node in the citation network. Finally, we removed the node representing the original literature review, leaving 1-hop, 2-hop, and 3-hop citation networks for each review. The mutual citations among references form complex citation networks, averaging 6,658.4 papers and 11,632.9 edges, including isolated papers.
A.2. Implementation
All experiments were conducted on a Linux-based server equipped with 4 NVIDIA A10G GPUs. For 518 review papers, we successfully collected taxonomy trees with hierarchical citation graph clustering labels for 313 of the literature reviews. Of these, 200 reviews were used to train the hierarchical citation graph clustering and taxonomy generation module, while the remaining 118 were used to test the performance of the pre-trained hierarchical citation graph clustering model. 318 reviews were reserved for a comprehensive evaluation of review content generation. The number of articles retrieved in retrieval phase was set to 200. The scaling factor $\alpha$ is set to 1.
Pre-Train Hierarchical Citation Graph Clustering Module. GNN used in this paper is GAT (Veličković et al., 2017) which has 2 layers with 4 heads per layer and a hidden dimension size of 1024. MLPϕ has 2 layers and a hidden dimension size of 1024. The edge connection threshold $p_{\tau}$ is searched in [0.1, 0.2, 0.5, 0.8]. The clustering model is trained for a maximum of 500 epochs using an early stop scheme with patience of 10. The learning rate is set to 0.001. The training batch is set to 512 and the test batch is 1024.
@@ A.2. Implementation
A.3. Evaluation Metrics
We engaged two groups of human raters, each consisting of two professional PhD students in computer science, to evaluate the generated taxonomies and literature reviews using two scoring mechanisms. Each group scores 100 generated samples Given the need to compare model-generated outputs with those written by humans, we designed the evaluation criteria to be straightforward and focused. The two key metrics used are Adequacy and Validity:
Adequacy is a binary metric where evaluators respond with “Yes” or “No” to the question: “Compared to the taxonomy written by humans, is this taxonomy suitable for learning this field?” This assesses whether the taxonomy is practically usable and meets the fundamental requirements for understanding the domain.
Validity, on the other hand, is rated on a scale from 1 to 5, evaluating the degree to which the taxonomy accurately reflects factual information and represents the domain’s conceptual structure. The scoring is defined as follows:
  • 1 – Completely inaccurate, with significant factual errors or misrepresentations of the domain.
  • 2 – Mostly inaccurate, capturing only a few correct facts but failing to represent the domain coherently.
  • 3 – Moderately accurate, containing some factual correctness but missing important concepts or relationships.
  • 4 – Mostly accurate, representing the domain well with minor factual inaccuracies or omissions.
  • 5 – Highly accurate, thoroughly reflecting the domain’s factual structure with no noticeable errors.
This combination of metrics allows us to capture both the practical usability and the factual correctness of the generated taxonomies, ensuring a comprehensive and nuanced evaluation from multiple perspectives.
For LLM evaluation, we assess the generated content from three perspectives: coverage, relevance, and structure, with each scored on a scale from 1 to 100. The specific prompts used for these evaluations are shown in Figure 5, Figure 6 and Figure 7.
Figure 5. Prompt used for evaluating coverage with LLMs.
Figure 6. Prompt used for evaluating relevance with LLMs.
Figure 7. Prompt used for evaluating structure with LLMs.
A.4. Investigation of Retrieval Models
We experimented with different retrieval models and strategies, testing two representative methods: the sparse retrieval model, BM25 (Robertson et al., 2009), and the dense retrieval model, SentenceBert (Reimers, 2019). In citation networks, neighbor information and the topological structure play a crucial role in retrieval, as papers on the same topic often cite each other. To assess the impact of using neighbor information, we applied two retrieval strategies for both models: one incorporating neighbor information as described in Section LABEL:sec:retrieval (Retrieval w/ Neighbor) and the other excluding neighbor information (Retrieval w/o Neighbor). Given a topic (specifically, the title of a review paper), we retrieved papers related to this topic from the citation network and measured the accuracy by calculating how many of the retrieved papers appeared in the references of the corresponding literature review. The number of retrieved papers was not fixed, but matched the reference count for each review paper.
Table 5. Results of retrieval on the citation network corresponding to 50 review papers. 2-hop and 3-hop represent citation networks of review papers at different scales. 1-hop (merged) refers to the 1-hop citation network of a review paper, merged with all other 1-hop citation networks, different review papers. Similarly, 2-hop (merged) is constructed by merging the 2-hop citation network of a review with all other 49 review papers.
@@ A.4. Investigation of Retrieval Models
A.5. The Choice of GNN
In addition to GAT (Veličković et al., 2017), we also explored other GNNs as graph encoders, i.e., GCN (Kipf and Welling, 2016) and Graph Transformer (Shi et al., 2020). The comparison results of these models on clustering are shown in Table 6.
Table 6. Performance of different GNN on hierarchical citation graph clustering.
@@ A.5. The Choice of GNN
Appendix B Literature Review Generation Example